OpenStudy (savannaho489):

In N2 + 3H2 → 2NH3, what is the actual yield if the theoretical yield of NH3 is 326 g and the percent yield is 85%? help and please explain!

OpenStudy (ciarán95):

\[Welcome~To~OpenStudy!!~:)\] Firstly, we'll define what the formula for the percentage yield is: \[Percentage~Yield = \frac{ Actual~Yield }{ Theoretical~Yield } \times 100\] \[Actual~Yield = \frac{ Percentage~Yield \times Theoretical~Yield }{ 100 }\] As far as I can see, we don't need to use the balanced equation for the reaction, as the numbers we need are already given to us in the question. We're told that the theoretical yield of our product, NH3, is 328g and that we have a percentage yield of 85%. So.... \[Actual~Yield~(g) = \frac{ (85) \times (326g) }{ 100 }\] which if you work out should give you the correct answer. Hope that helps! :)